p,q,r are three positive numbers and Q=(p+q+r)÷2. If (Q-p):(Q-q):(Q-r)=2:5:7. Then find the ratio of p:q:r.

Grade:10

1 Answers

Arun
6 years ago
 

Dear student

 

Q-p/Q-q/Q-r = 2:5:7

 

Q-p = 2K                  ........1

 

Q-q = 5k                  ..........2

 

Q-r = 7k                 .........3

 

adding all these eq

 

3Q -(p+q+r) = 14k         ...........4

 

p+q+r/2 = Q              (given)

 

using this eq 4 becomes

 

Q = 14k

 

putting Q in 1 ,2 ,3

 

p = 12k , q = 9k , r = 7k

 

ratio = p:q:r = 12:9:7

 

Regards

Arun (askIITians forum expert)